OpenStudy (samigupta8):

If g(x)=max|y^2-xy| then the minimum value of g(x) for real x is.... Here y can go from [0,1]

OpenStudy (samigupta8):

@mayankdevnani @parthkohli @priyar..help

OpenStudy (sparrow2):

is here absolute value? if so then minimum value is 0

OpenStudy (samigupta8):

Lol...bt not in the options....nvm

OpenStudy (samigupta8):

N yes it's definitely an absolute value...

OpenStudy (samigupta8):

@imqwerty WUd u help

imqwerty (imqwerty):

okay 1st we gotta find the stationary point by taking \(f(y)=y^2-xy\) and solving \(\large \frac{df}{dy}=0\) we will get \(y=\large \frac{x}{2}\) now this can be local minimum or local maximum so we gotta compare it with extremities \(f(0)\) and \(f(1)\) \(\begin{align} g(x) &= \max\left(f(0)\;,f(1),f\left(\frac{x}{2}\right)\right)\\ &= \max\left(0\;,1-x,\frac{-x^2}{4}\right)\\ &= \max\left(0\;,1-x\right)\\ &= \begin{cases} 0 & x>1 \\ 1-x & x\le 1 \end{cases} \end{align}\)

OpenStudy (samigupta8):

Correct then....

OpenStudy (samigupta8):

As PR the graph we can get its minimum value as 0 from ur method....

OpenStudy (samigupta8):

N actually qwerty that was the function in x how could u differentiate it vid respect to x ?

OpenStudy (samigupta8):

Ayush btao Naa

imqwerty (imqwerty):

sry sry i didn't see ur cmnts here

OpenStudy (samigupta8):

So see it now....lol

imqwerty (imqwerty):

yes \:) Note that \(g(x)\) is \(\ge 0\) and \(x \le 1\). so the minimum value of \(g(x)\) is clearly 0.

OpenStudy (mayankdevnani):

what are the options then ?

OpenStudy (samigupta8):

1/4 1/2 3+√8 3-√8

OpenStudy (dumbcow):

continuing @imqwerty response \[g(x) = \max(0, |1-x|, |\frac{-x^2}{4}|)\] If we graph these we can find the point where g(x) is smallest  \[1-x = \frac{x^2}{4} \rightarrow x^2 +4x - 4 = 0\] \[x = 2 \sqrt{2} - 2\] \[\min(g(x)) = 1-x = 1 - (2\sqrt{2}-2) = 3 - 2 \sqrt{2}\]